# On the Costs and Benefits of Learned Indexing for Dynamic High-Dimensional Data: Extended Version

### v1

@@ Abstract
One of the main challenges within the growing research area of learned indexing is the lack of adaptability to dynamically expanding datasets. This paper explores the dynamization of a static learned index for complex data through operations such as node splitting and broadening, enabling efficient adaptation to new data. Furthermore, we evaluate the trade-offs between static and dynamic approaches by introducing an amortized cost model to assess query performance in tandem with the build costs of the index structure, enabling experimental determination of when a dynamic learned index outperforms its static counterpart. We apply the dynamization method to a static learned index and demonstrate that its superior scaling quickly surpasses the static implementation in terms of overall costs as the database grows. This is an extended version of the paper presented at DAWAK 2025.
Keywords:
Learned indexing Dynamization Dynamic datasets k-NN search ANN search
@@ 1 Introduction
This paper is focused on the latter approach, proposing a dynamization method capable of transforming a static learned indexing structure for complex, high-dimensional data into a dynamic one. This is achieved through generalized node-splitting and node-broadening operations, along with a set of simple rules for their application. Even though it was developed with a learned indexing use case in mind, this approach is general enough to be applicable to any partition-based indexing technique that struggles with dynamization.
@@ 2.1 Learned Indexing
Learned indexing has emerged as a novel approach to indexing, based on the premise that indexes can be viewed as models [15]. While there is a growing body of research on using such indexes for one-dimensional datasets [32], the straightforward idea of learning the cumulative distribution function is not directly applicable to high-dimensional data. Most of the existing learned indexes for high-dimensional data either create their own clustering [8, 31, 13, 18] or learn an existing one [1, 17]. The limitation of such approaches is that they are designed to index static datasets only. They do not address the dynamic indexing scenario, aside from resorting to a complete rebuild of the index [17] without any guiding policy, essentially relying on the user to trigger the reconstruction process manually.
@@ 2.1.1 Insert Operation
After a certain point, the accuracy of a learned index will inevitably deteriorate due to the continuous insertion of new objects that exhibit distribution shifts, or deletion of existing objects. Adapting a trained model to new data requires fine-tuning, which can overwrite parameters critical for handling previously indexed data. In neural networks, this phenomenon is known as catastrophic forgetting [24, 14], where optimizing for new tasks degrades performance on earlier ones. A subfield of machine learning called continual learning [35] studies methods that allow models to adapt to situations where the learned knowledge is to be adjusted throughout the model’s existence. Such methods attempt to balance the trade-off between learning plasticity, i.e., how well the new task is learned, and memory stability, i.e., how well the model remembers previous knowledge.
@@ 2.2.2 Graph-based
Graph-based indexes [36] are considered the state of the art in approximate nearest neighbor search [2]. Conceptualized as proximity graphs [11, 23], they approximate their theoretical counterparts [10, 34], reducing building costs and accelerating search by omitting non-relevant edges. HNSW [22], the most researched index in this category, is incorporated into a wide range of vector databases [27] and is dynamic by design due to its incremental building procedure. However, it lacks a hard delete operation, instead relying on so-called lazy deletion, which depends on the eventual garbage collection of tagged objects. [^1] https://github.com/nmslib/hnswlib/issues/4
@@ 2.2.3 LSH-based
Most LSH-based methods [12, 21, 19] are built for static use and thus fail to efficiently accommodate dynamic data loads. Adding new data to a created LSH typically requires recalculating the hash functions and updating the index. Two notable works attempt to extend the static nature of LSH to dynamic use cases: LSH forest [3] through the use of a multiple prefix tree structure and PLSH [33] by employing an insert-optimized hash table.
@@ 3.1 Dynamized Learned Index
Input: structure $S$, a data node $n$, # of new children $n\_child$, distance function $d$
@@
Input: structure $S$, an inner node $n$, # of new children $n\_child$, distance function $d$
@@
The shorten operation addresses the removal of severely underpopulated nodes. This involves deleting the node and reinserting its objects into the index by removing the corresponding output neuron and its connections from the learned model (MLP) – see Algorithm 3 and Figure 3. Unlike adding a neuron (which requires global retraining and risks catastrophic forgetting), this localized operation allows well-populated categories to redistribute the obsolete category’s data points, eliminating its decision boundary. While effective for shallower MLPs and underpopulated categories, deeper networks or significant data distribution shifts may necessitate fine-tuning to maintain model performance.
Policies that trigger the structural changes. After defining the three basic operations for a structural update (deepen, broaden, and shorten), the next step is to establish policies to invoke them. First, we define minimum and maximum bounds on the leaf node capacity to impose limits on the costs of sequential search. Similarly, these bounds are defined on the inner nodes for optimal discriminative power of a single model. We implement a policy of detecting and resolving any violations of these bounds – in case of a leaf node having fewer than 5 objects (underflow), we invoke the shorten operation. To manage overflow, the structure ensures that the average occupancy of leaf nodes remains below 1 000. When this bound is violated, we invoke node extension, alternating between deepen and broaden operations to maintain a shallow index. Specifically, a node will be deepened (split) until the index reaches a maximum depth of two levels. After this depth is reached, the broaden operation is invoked on the appropriate node to accommodate further growth, re-partitioning, and retraining of the learned model.
@@
S.insert($A$); S.delete_nodes($arr_{n}$); return S;
@@ 3.3 Amortized Cost Model
An ideal evaluation metric should combine the search costs and build costs into a single objective that can be compared on a per-query basis regardless of how dynamic the method is. A previous work evaluated the total cost of executing a fixed number of queries (including the build costs) [37] – however, this approach assumes a read-only use case, with all the data provided upfront rather than being inserted dynamically as the queries arrive. While this metric can be appropriate for read-dominated workloads, it may not fully capture the performance of systems that are subject to frequent updates or dynamic changes.
To explore how such systems perform in more dynamic settings, let us define a new property of the system: the relative number of queries per inserted object. We will call this the querying frequency, and define it simply as #$\frac{\#queries}{\#new\_objects}$. We will assume querying frequency to be an inherent property of the given indexing scenario, which does not change during the lifetime of the database. This allows us to relate the number of new objects added to the database (which dictate the rebuild costs of a dynamic structure) with the number of queries performed on the database (which determine the amortization).
@@
In other words, if an index is rebuilt after adding 1K new objects, and the querying frequency is 100 queries per new object, then one build of the index will last for 100K queries. Therefore, the amortized cost of a query should include $\frac{1}{100000}$ of the index build costs in addition to the immediate search costs.
@@ 4 Experiments
For our experimental evaluation, we use SIFT descriptors [20], one of the datasets used in ANN-benchmarks [2]. This dataset consists of 1 million objects of 128 dimensions with the Euclidean distance metric, and additionally contains 10K queries in a 30-NN setup for experimental evaluation.
The experiments were conducted using Python 3.8 with PyTorch 1.8 on a machine with AMD EPYC 7261 8-Core Processor and 50 GiB of RAM. GPU acceleration was not used.
@@
As described in Section 3, we compare the dynamized version of the learned index against two baselines – the Naive rebuild baseline, which rebuilds the entire index periodically based on a pre-defined rebuild interval, and the No rebuild baseline, which simply stores incoming objects without adjusting its structure until it runs out of experimental data to process. We performed the experiment with various database sizes, ranging from 100K to 900K. For the baselines, this means that the database was first built using the given number of initial objects, and then new objects were added – in the case of the Naive rebuild baseline, the objects were added until a rebuild was triggered, in the No rebuild baseline, they were added until all 1M experimental data objects were indexed. The dynamized index was built gradually starting with no objects, and its amortized performance was simply evaluated after every 100K objects as the index grew and adapted.
Since the Naive rebuild baseline is sensitive to the selection of rebuild interval, we optimized this parameter for each of the 4 experimental scenarios, resulting in 4 different performance curves of this method. As a result, in each evaluated scenario, one of the Naive rebuild baselines shows the optimal solution (since the rebuild frequency is set up perfectly for that scenario), and the other Naive rebuild baselines show the penalty for parameterizing the method incorrectly.
As for the No rebuild baseline, since the index is only built once (using the initial number of objects), the build costs are as low as possible for the given initial database size, but the quality of queries will keep deteriorating towards exhaustive search in the limit. Due to the size of our experimental dataset (1M objects), we can only simulate smaller and smaller increments of this deterioration as we use more objects to build the initial database. In other words, when the initial database size is 100K, the plotted results show how much queries deteriorate if database size increases by 900% with no rebuild; when we start with 900K objects, the plot shows how much performance deteriorates if we add $\sim$11additional objects, as there are no more objects to insert. Thus, when plotting the amortized cost, the results of this baseline keep converging towards the dynamized methods, not due to improved performance of the method, but because the fixed size of the experimental dataset limits the ability to model larger-scale growth.
@@
Note that the static index used in the Naive rebuild and No rebuild baselines is a single-level structure, implemented as a single MLP, and is parameterized to hold an average of 1K objects per bucket. Additionally, note that only insertion of new objects is considered – deletion and re-balancing can also be implemented in a straightforward manner (as shown in Section 3.1), but the experimental evaluation of such operations is beyond the scope of this paper.

### v2

# On the Costs and Benefits of Dynamizing a Learned Index

@@ Abstract
One of the main challenges within the growing research area of learned indexing is the lack of adaptability to dynamically expanding datasets. This paper explores the dynamization of a static learned index for complex data through operations such as node splitting and broadening, enabling efficient adaptation to new data. Furthermore, we evaluate the trade-offs between static and dynamic approaches by introducing an amortized cost model to assess query performance in tandem with the build costs of the index structure, enabling experimental determination of when a dynamic learned index outperforms its static counterpart. We apply the dynamization method to a static learned index and demonstrate that its superior scaling quickly surpasses the static implementation in terms of overall costs as the database grows.
@@ 1 Introduction
This paper is focused on the latter approach, proposing a dynamization method capable of transforming a static learned indexing structure into a dynamic one through generalized node-splitting and node-broadening operations, as well as a set of simple rules for their usage. Even though it was developed with a learned indexing use case in mind, this approach is general enough to be applicable to any partition-based indexing technique that struggles with dynamization.
@@ 2.1 Learned Indexing
Learned indexing [16, 33] has emerged as one of the novel approaches to indexing, based on the premise that indexes can be viewed as models [16]. However, the straightforward idea of learning cumulative distribution function for one-dimensional data is not directly applicable to high-dimensional data. Most of the existing learned indexes for high-dimensional data either create their own clustering [8, 32, 13, 19] or learn an existing one [1, 18]. The limitation of such approaches is that they are designed to index static datasets only. They do not address the dynamic indexing scenario, aside from resorting to a complete rebuild of the index [18] without any guiding policy, essentially relying on the user to trigger the reconstruction process manually.
@@ 2.1.1 Insert Operation
After a certain point, the accuracy of a learned index will inevitably deteriorate due to the continuous insertion of new objects that exhibit distribution shifts, or deletion of existing objects. Furthermore, adapting a trained model within a learned index to a new distribution results in a phenomenon called catastrophic forgetting [25, 15] in which the model forgets the already learned knowledge. A subfield of machine learning called continual learning [36] studies methods that allow models to adapt to situations where the learned knowledge is to be adjusted throughout the model’s existence. Such methods attempt to balance the trade-off between learning plasticity, i.e., how well the new task is learned, and memory stability, i.e., how well the model remembers previous knowledge.
@@ 2.2.2 Graph-based
Graph-based indexes [37] are considered the state of the art in approximate nearest neighbor search [2]. Conceptualized as proximity graphs [11, 24], they approximate their theoretical counterparts [10, 35], reducing building costs and accelerating search by omitting non-relevant edges. HNSW [23], the most researched index of this category, is incorporated into a wide range of vector databases [28] and is dynamic by design due to its incremental building procedure, although without a delete operation [^1] https://github.com/nmslib/hnswlib/issues/4.
@@ 2.2.3 LSH-based
Most LSH-based methods [14, 12, 22, 20] are built for static use and thus fail to efficiently accommodate dynamic data loads. Adding new data to a created LSH typically requires recalculating the hash functions and updating the index. Two notable works attempt to extend the static nature of LSH to dynamic use cases: LSH forest [3] through the use of a multiple prefix tree structure and PLSH [34] by employing an insert-optimized hash table.
@@ 3.1 Dynamized Learned Index
Input: structure $S$, a data node $n$, # of new children n_childn\_child, distance function $d$
@@
Input: structure $S$, an inner node $n$, # of new children n_childn\_child, distance function $d$
@@
Removal of a defective or obsolete node. In case of a severe underflow of a data node or other type of defect, this node should be deleted and its objects (if any) reinserted using the shorten operation. This operation is implemented by removing a category from among the leaf nodes, which in our case involves removing one of the learned model’s output neurons. Unlike the insertion of an additional neuron, this removal does not affect the model’s prediction mechanism. For more details, see Algorithm 3 and Figure 3.
Policies that trigger the structural changes. After defining the basic operations for a structural update, the next step is to establish policies to invoke them. First, it is reasonable to define minimum and maximum bounds on the leaf node capacity to impose limits on the costs of sequential search. Similarly, such bounds should be defined on the inner nodes for optimal discriminative power of a single model. We chose to define a policy of detecting and resolving any violations of these bounds – in case of a leaf node having less than 5 objects (underflow), we immediately delete it and redistribute its objects among the other nodes. To manage overflow, the structure ensures that the average occupancy of all leaf nodes remains below 1 000. When this bound is violated, nodes with the highest occupancy are extended until the average falls back below the threshold. In terms of depth, shallow indexes are preferred – for the 1 000 000 objects in our experimental dataset, it only grows up to a maximum of two levels.
@@
S.insert(AA); S.delete_nodes($arr_{n}$); return S;
@@ 3.3 Amortized Cost Model
An ideal evaluation metric should combine the search costs and build costs into a single objective that can be compared on a per-query basis regardless of how dynamic the method is. Thus, in addition to the search costs of a query, which are easily determined, we should also consider the amortized build costs of that query – that is, the total build costs of a structure divided by the number of queries performed on that structure.
However, the number of queries can only be determined once we define a new property of the system: the relative number of queries per inserted object. We will call this the querying frequency, and define it simply as $\frac{\#queries}{\#new\_objects}$. We will assume querying frequency to be an inherent property of the given indexing scenario, which does not change during the lifetime of the database. This allows us to relate the number of new objects added to the database (which dictate the rebuild costs of a dynamic structure) with the number of queries performed on the database (which determine the amortization).
@@
In other words, if an index is rebuilt after adding 1 000 new objects, and the querying frequency is 100 queries per new object, then one build of the index will last for 100 000 queries. Therefore, the amortized cost of a query should include $\frac{1}{100000}$ of the index build costs in addition to the immediate search costs.
@@ 3.3.1 Optimal Rebuild Interval
For additional details on how the optimal rebuild interval varies based on the database size and indexing scenario, and the relative influence of search versus build costs, please refer to Tables 1 and 2 in the Appendix.
@@ 4 Experiments
For our experimental evaluation, we use SIFT descriptors [21], one of the datasets used in ANN-benchmarks [2]. This dataset consists of 1 million objects of 128 dimensions with the Euclidean distance metric, and additionally contains 10 000 queries in a 30-NN setup for experimental evaluation.
@@
As described in Section 3, we compare the dynamized version of the learned index against two baselines – the Naive rebuild baseline, which rebuilds the entire index periodically based on a pre-defined rebuild interval, and the No rebuild baseline, which simply stores incoming objects without adjusting its structure until it runs out of experimental data to process. We performed the experiment with various database sizes, ranging from 100 000 to 900 000. For the baselines, this means that the database was first built using the given number of initial objects, and then new objects were added – in the case of the Naive rebuild baseline, the objects were added until a rebuild was triggered, in the No rebuild baseline, they were added until all 1 000 000 experimental data objects were indexed. The dynamized index was built gradually starting with no objects, and its amortized performance was simply evaluated after every 100 000 objects as the index grew and adapted.
Since the Naive rebuild baseline is sensitive to the selection of rebuild interval, we optimized this parameter for each of the 4 experimental scenarios, resulting in 4 different performance curves of this method. As a result, in each evaluated scenario, one of the Naive rebuild baselines shows the optimal solution (since the rebuild frequency is set up perfectly for that particular scenario), and the other Naive rebuild baselines show the penalty for parameterizing the method incorrectly.
As for the No rebuild baseline, since the index is only built once (using the initial number of objects), the build costs are as low as possible for the given initial database size, but the quality of queries will keep deteriorating towards exhaustive search in the limit. Due to the size of our experimental dataset (1 000 000 objects), we can only simulate smaller and smaller increments of this deterioration as we use more objects to build the initial database. In other words, when the initial database size is 100 000, the plotted results show how much queries deteriorate if database size increases by 900% with no rebuild; when we start with 900 000 objects, the plot shows how much performance deteriorates if we add $\sim$11% additional objects, as there are no more objects to insert. Thus, when plotting the amortized cost, the results of this baseline keep converging towards the dynamized methods, not due to improved performance of the method, but because the fixed size of the experimental dataset limits the ability to model larger-scale growth.
Note that the static index used in the Naive rebuild and No rebuild baselines is a single-level structure, implemented as a single MLP, and is parameterized to hold an average of 1 000 objects per bucket. Additionally, note that only insertion of new objects is considered – deletion and re-balancing can also be implemented in a straightforward manner (as shown in Section 3.1), but the experimental evaluation of such operations is beyond the scope of this paper.
@@
5 Acknowledgments
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic (ELIXIR-CZ, Grant No. LM2023055), e-INFRA CZ (ID:90254), and the Czech Science Foundation [GF23-07040K].
@@ 6 In Conclusion
Appendix
The following tables detail the Naive rebuild method parameterized for various indexing scenarios. Table 1 shows the optimal number of new objects that should trigger a full rebuild of the index. As a reminder, QPI refers to the number of queries per insert and TR refers to target recall in a given scenario.
The rebuilds are most frequent in a scenario with high querying intensity and high target recall, since this scenario is most sensitive to deterioration in search quality – it is hard to reach high recall in a decayed index, and since there are many queries, the penalty for slow searching quickly outpaces the costs of rebuilding the index from scratch. The opposite is true for the low querying intensity—low target recall scenario, where the penalty for deteriorating search quality is relatively mild – it is still feasible to reach a recall of 0.5 in a low-quality structure, and since there are few queries, it takes a long time for the slower search to incur so much overhead that a full rebuild becomes worthwhile.
The single N/A value in each table is due to the limited number of objects in the dataset – if the initial database size is 900 000 but the rebuild interval would be higher than 100 000, we run out of objects before even a single index rebuild could be triggered and evaluated.
DB size 100 QPI 100 QPI 1 QPI 1 QPI
TR 0.9 TR 0.5 TR 0.9 TR 0.5
100 000 2 700 7 900 29 900 70 000
200 000 4 300 11 500 30 000 110 000
300 000 4 800 16 500 30 000 120 000
400 000 6 000 20 600 40 000 160 000
500 000 5 900 23 600 40 000 180 000
600 000 6 000 29 700 40 000 190 000
700 000 8 000 30 000 50 000 240 000
800 000 10 600 30 000 50 000 190 000
900 000 6 600 30 000 60 000 N/A
Table 1: The optimal rebuild intervals for the Naive rebuild method in different indexing scenarios.
DB size 100 QPI 100 QPI 1 QPI 1 QPI
TR 0.9 TR 0.5 TR 0.9 TR 0.5
100 000 98.53% 97.23% 89.29% 84.77%
200 000 99.08% 97.30% 96.46% 88.75%
300 000 99.19% 97.75% 96.59% 89.06%
400 000 99.34% 97.97% 97.45% 91.10%
500 000 99.32% 98.01% 97.76% 92.37%
600 000 99.36% 98.33% 97.80% 92.35%
700 000 99.51% 99.19% 98.20% 94.04%
800 000 99.63% 99.14% 98.27% 92.18%
900 000 99.38% 99.06% 98.55% N/A
Table 2: The percentage of search costs within the overall amortized costs per query.